# Robust Spatial Confounding Adjustment via Basis Voting

### v1

@@ 6 Real Data Analysis
We illustrate the robustness of our mode estimator in an analysis studying the effect of air quality on COVID-19 mortality. We utilize the county-level COVID-19 dataset analyzed in Guan et al. (2023) and originally collected by Wu et al. (2020), which includes comprehensive information for $n=3109$ U.S. counties. For each county, the outcome is the cumulative number of COVID-19 deaths through May 12, 2020 (originally obtained from Johns Hopkins University Coronavirus Resource Center), scaled by the county’s population. The primary exposure variable is the long-term average PM2.5 concentration, computed by averaging estimates from an established exposure prediction model over the years 2000–2016. Figure 7 displays the scaled death counts and PM2.5 concentrations for each county. Counties with zero death counts are shown in grey in the figure.
@@
In addition to the main exposure and outcome variables, the data set contains 20 additional county-level covariates which are potential confounders. To transform this data into a standard geostatistical one, we assign each county the coordinates of its geographic centroid. Counties for which the boundary centroids could not be calculated were dropped from the analysis resulting in an effective sample size of $n=3082$. Consistent with our focus on linear models, we apply a logarithmic transformation to the scaled mortality outcome, replacing zeros with a small positive value to avoid loss of observations and ensure stability of the log transformation. The ordinary least square estimate of the mortality rate ratio for 1 unit increase in PM2.5 even after adjusting for the measured confounders is $\exp(-1.286)=0.267$ which shows a counterintuitive result that higher pollutant levels are protective against COVID-19 mortality. This paradoxical finding is also discussed in Wu et al. (2020) and Guan et al. (2023), and we agree with their conclusion that there is a possibility for the presence of unmeasured spatial confounding.
According to Guan et al. (2023), after adjusting for spatial confounding, the estimated percentage increase in mortality rate associated with a one-unit increase in PM2.5 is approximately 16%, i.e., the effect size is around $1.16$. To assess the robustness of our proposed approach relative to existing methods that assume an ordered set of basis functions for adjustment, we consider two classes of spatial basis functions: (1) the conventional thin-plate regression splines (TPRS) orthogonalised over the set of locations which exhibit an inherent ordering of smoothness, and (2) the cartesian product of two one-dimensional Fourier bases, which lack such canonical ordering. As a benchmark, we compare our mode-based estimator to the standard basis adjustment approach in regression, which is known to yield consistent results when basis functions are ordered by increasing roughness.
Figure 8: Results for the COVID-19 Data: Estimated mortality-rate ratio for an increase of $1\mu g/m^{3}$ increase in PM2.5 for different numbers of basis functions. The results are for two sets of basis functions: (a) when the basis functions have a canonical ordering of increasing roughness, and (b) when the basis functions do not have a canonical ordering among the basis functions. The blue line is for the model adjusted for $k$ basis functions along with the measured covariates, and the red is for basis voting, our mode-based estimator. Black dotted line is the estimate of 1.16 or 16% reported by Guan et al. (2023).
@@
As anticipated, when the basis set has an ordering of smoothness (as with TPRS), the results of both estimators are highly similar and align with the estimate from the previous analyses. Notably, when utilizing the unordered Fourier basis, our mode-based estimator continues to produce estimates close to the presumed effect, whereas the standard basis adjustment method fails to control for unmeasured spatial confounding, producing a very different estimate. This result underscores the advantage of our approach in settings where the exposure does not represent a ‘rougher’ function of space than the confounder in terms of a given basis expansion.
@@ 7 Conclusion and Discussion
Unmeasured spatial confounding poses a major challenge for effect size estimation in spatial linear models, particularly when exposures and confounders share smooth spatial structure. Existing approaches often assume exposures vary more roughly than confounders, but this assumption breaks down in many environmental health settings where exposures themselves are smooth or when using basis functions where there is no single notion of smoothness.
We address this gap by modeling exposures and confounders through basis expansions and emphasizing the fact that a sufficiently large collection of basis functions contributing to the differential support of the exposure is sufficient for identifying the causal effect. Some of the basis functions act as valid candidates, while others induce confounding. Leveraging this structure, we introduce basis voting, a plurality-rule based process to arrive at an estimator that identifies causal effects without requiring prior knowledge of valid bases or bounds on invalid ones. We provide theoretical guarantees for consistency and propose a provably more efficient candidate estimator with improved asymptotic variance. Simulations and a real data analysis confirm that our method delivers unbiased estimates whenever exposure and confounder signals are separable on a plurality of basis functions, regardless of their relative smoothnesses.
This work advances spatial causal inference by relaxing restrictive smoothness assumptions and unifying existing plurality-based strategies under a robust IV perspective. While our framework is limited to linear effects, we plan to explore future extensions to nonlinear models. Also, developing valid statistical inferential procedures using the mode-estimator is an important future direction. Overall, plurality-based estimation provides a flexible and robust tool for addressing spatial confounding in complex applied settings.
@@ 8 Acknowledgement
The work was partly supported by National Institute of Health Sciences grant R01 ES033739 and ONR grant N000142412701. The authors acknowledge the use of ChatGPT.com for improving the exposition in some parts and for some assistance with developing the theoretical results.
@@
S1 Proofs
S1.1 Notations
@@
We denote $\mathbb{R}$ and $\mathbb{N}$ to be the set of real numbers and natural numbers respectively. For a sequence of random variables, $\xrightarrow[n\to\infty]{P}$ denotes convergence in probability and $\xrightarrow[n\to\infty]{d}$ denotes weak convergence. We will also use stochastic ordering notations throughout the proof. For a sequence of random variable $\{X_{n}\}$ and a sequence of real numbers $\{a_{n}\}$, we say $X_{n}=\mathcal{O}_{P}(a_{n})$ if for all $\epsilon>0$ there exists a $M>0$, such that for all $n\in\mathbb{N}$, we have $\mathbb{P}\left(\left|\frac{X_{n}}{a_{n}}\right|>M\right)<\epsilon$. Also, we say $X_{n}=o_{P}(a_{n})$ if for all $\epsilon>0$, $\mathbb{P}\left(\left|\frac{X_{n}}{a_{n}}\right|>\epsilon\right)\xrightarrow[]{n\to\infty}0$. For a real valued function $f$ defined on an inner-product space $(\mathcal{M},\langle\cdot,\cdot\rangle)$, we denote $\|f(\cdot)\|_{\mathcal{M}}\coloneqq\sup_{t\in\mathcal{M}}\langle f(t),f(t)\rangle$. All the integrals are over the support $\mathcal{D}$ of the sampling density $f_{S}$, and hence unless specified, by $\int$ we will mean $\int_{\mathcal{D}}$.
S1.2 Proof of Lemma 1
Proof.
and the uniqueness follows from $n_{t}>M$.
$\varepsilon\;<\;\frac{K(0)}{2m}\gamma.$ (8)
and thus
Evaluating at $x=t$ gives
[…]
By the choice (8), the right-hand side is strictly positive.
hence for any $x\in A$ and $0<h\leq h_{0}$,
Now fix any $r>0$, since
$KD(\bm{a}_{m},K,h,t)>KD(\bm{a}_{m},K,h,x),$
Hence $t$ is the unique global mode in this setup. ∎
S1.3 Proof of Theorem 1
Proof.
For the proof of this corollary, we will use a result from Vaart & Wellner (1997) on the convergence of argmax of stochastic processes.
Lemma S1.1.
@@
[…]
$\|KD(\mathcal{C}_{n},K,h,x)-KD(\beta+R_{d},X,h,x)\|_{\mathbb{R}}$
$=$ $\underset{x\in\mathbb{R}}{\sup}\left|KD(\mathcal{C}_{n},K,x)-KD(\beta+R_{d},K,x)\right|$
$=$ $\underset{x\in\mathbb{R}}{\sup}\left|\dfrac{1}{mh}\sum_{k\in S_{x,d}}K\left(\dfrac{x-\hat{\beta}_{k,n}}{h}\right)-\dfrac{1}{mh}\sum_{k\in S_{x,d}}K\left(\dfrac{x-(\beta+r_{k,d})}{h}\right)\right|$
$\leq$ $\underset{x\in\mathbb{R}}{\sup}\dfrac{1}{mh}\sum_{k\in S_{x,d}}\left|K\left(\dfrac{x-\hat{\beta}_{k,n}}{h}\right)-K\left(\dfrac{x-(\beta+r_{k,d})}{h}\right)\right|$
$\leq$ $\underset{x\in\mathbb{R}}{\sup}\dfrac{L_{K}}{mh}\sum_{k\in S_{x,d}}\left|\left(\dfrac{x-\hat{\beta}_{k,n}}{h}\right)-\left(\dfrac{x-(\beta+r_{k,d})}{h}\right)\right|\quad\text{(using Lipschitz Continuity)}$
$=$ $\dfrac{L_{K}}{mh^{2}}\sum_{k\in S_{x,d}}\left|\hat{\beta}_{k,n}-(\beta+r_{k,d})\right|\xrightarrow[n\to\infty]{P}0$
Now, based on the Plurality Rule Assumption 2, and Lemma 1, it is ensured that for $\theta_{0}=\beta$, we have $KD(\beta,K,h,\theta_{0})>\underset{\theta\notin G}{\sup}KD(\theta+R_{d},K,h,\theta)$ for any open set $G$ that contains $\theta_{0}$.
Thus defining $\hat{\beta}_{n}=\underset{x}{\operatorname*{arg\,max}}\ KD(\mathcal{C}_{n},K,h,x)$, we can say that $KD\left(\mathcal{C}_{n},K,h,\hat{\beta}_{n}\right)\geq\sup_{\theta}KD\left(\mathcal{C}_{n},K,h,\theta\right)-o_{P}(1)$. Thus, using Lemma S1.1, we can say that $\hat{\beta}_{n}\xrightarrow[n\to\infty]{P}\beta$. ∎
Before starting the proof for the next theorems, we will state a few Lemmas which will have the building blocks that we will be needing for the proof:
@@
S1.4 Proof of Theorem 2
@@
Recall that $\bm{H}_{1:d}=\begin{bmatrix}\bm{h}_{n,1}&\cdots&\bm{h}_{n,d}\end{bmatrix}$ and $\bm{P}_{d}$ is the projection matrix on its column space. Let us choose some $j\in S_{x,d}$, which means $\alpha_{x}^{(j)}\neq 0$. Now, let us look at $\hat{\beta}_{j,n}^{PROJ}$:
@@
$\leq$ $\underset{u\geq 1}{\sup}\ \alpha_{x}^{(u)}\sum_{u=1}^{k}\mathbb{I}(u=j)\leq\underset{u\geq 1}{\sup}\ \alpha_{x}^{(u)}\leq\sum_{u\geq 1}\alpha_{x}^{(u)^{2}}<\infty$
@@
Similarly we have $\frac{1}{n}\bm{h}_{n,j}^{\top}\bm{U}_{n}\xrightarrow[n\to\infty]{P}\alpha_{u}^{(j)}$. Lemma S1.2 implies that $\frac{1}{n}\bm{h}_{n,j}^{\top}\bm{\epsilon}_{n}\xrightarrow[n\to\infty]{P}0$. Thus for $j\in S_{x,d}$ i.e. when $\alpha_{x}^{(j)}\neq 0$, we have $\hat{\beta}_{j,n}^{PROJ}\xrightarrow[n\to\infty]{P}\beta+\alpha_{u}^{(j)}/\alpha_{x}^{(j)}=\beta+r_{j,d}$ which completes the proof of consistency. Let us define $A_{i}=h_{j}(S_{i})[U(S_{i})+\epsilon_{i}]$ and $B_{n}=h_{j}(S_{i})X(S_{i})$,which means $\frac{1}{n}\bm{h}_{n,j}^{\top}[\bm{U}_{n}+\bm{\epsilon}_{n}]=\frac{1}{n}\sum_{i=1}^{n}h_{j}(S_{i})[U(S_{i})+\epsilon_{i}]=\bar{A}_{n}$ and $\frac{1}{n}\bm{h}_{n,j}^{\top}\bm{X}_{n}=\frac{1}{n}\sum_{i=1}^{n}h_{j}(S_{i})X(S_{i})=\bar{B}_{n}$. We will try to derive the asymptotic distribution of $\bar{A}_{n}/\bar{B}_{n}$, which is the bias of the estimator $\hat{\beta}^{PROJ}_{j,n}$. Now we have seen that $\mathbb{E}A_{i}=\mathbb{E}h_{j}(S_{i})[U(S_{i})+\epsilon_{i}]=\mathbb{E}h_{j}(S_{i})U(S_{i})=\int h_{j}(s)U(s)f_{S}(s)ds=\alpha_{u}^{(j)}$ and $\mathbb{E}B_{i}=\mathbb{E}h_{j}(S_{i})X(S_{i})=\alpha_{x}^{(j)}$ Also, notice that
@@
Since $\int(h_{j}U(s))^{2}f_{S}ds<\infty$ and $\int(h_{j}X(s))^{2}f_{S}ds<\infty$ as the functions are elements of $\mathcal{L}^{4}(\mathcal{D},f_{S})$, by Multivariate Central Limit Theorem (CLT), we have:
@@
$\sqrt{n}\left(\dfrac{\bar{A}_{n}}{\bar{B}_{n}}-\underbrace{\dfrac{\alpha_{u}^{(j)}}{\alpha_{x}^{(j)}}}_{r_{j,d}}\right)\xrightarrow[n\to\infty]{d}N\left(0,\sigma^{2}_{j,d}\right)$ (15)

### v2

@@ 6 Real Data Analysis
We illustrate the robustness of our mode estimator in an analysis studying the effect of air quality on COVID-19 mortality. We utilize the county-level COVID-19 dataset analyzed in guan2023spectral and originally collected by wu2020air, which includes comprehensive information for $n=3109$ U.S. counties. For each county, the outcome is the cumulative number of COVID-19 deaths through May 12, 2020 (originally obtained from Johns Hopkins University Coronavirus Resource Center), scaled by the county’s population. The primary exposure variable is the long-term average PM2.5 concentration, computed by averaging estimates from an established exposure prediction model over the years 2000–2016. Figure 7 displays the scaled death counts and PM2.5 concentrations for each county. Counties with zero death counts are shown in gray in the figure.
@@
In addition to the main exposure and outcome variables, the data set contains 20 additional county-level covariates which are potential confounders. To transform this data into a standard geostatistical one, we assign each county the coordinates of its geographic centroid. Counties for which the boundary centroids could not be calculated were dropped from the analysis resulting in an effective sample size of $n=3082$. Consistent with our focus on linear models, we apply a logarithmic transformation to the scaled mortality outcome, replacing zeros with a small positive value to avoid loss of observations and ensure stability of the log transformation. The ordinary least square estimate of the mortality rate ratio for 1 unit increase in PM2.5 even after adjusting for the measured confounders is $\exp(-1.286)=0.267$ which shows a counterintuitive result that higher pollutant levels are protective against COVID-19 mortality. This paradoxical finding is also discussed in wu2020air and guan2023spectral, and we agree with their conclusion that there is a possibility for the presence of unmeasured spatial confounding.
According to guan2023spectral, after adjusting for spatial confounding, the estimated percentage increase in mortality rate associated with a one-unit increase in PM2.5 is approximately 16%, i.e., the effect size is around $1.16$. To assess the robustness of our proposed approach relative to existing methods that assume an ordered set of basis functions for adjustment, we consider two classes of spatial basis functions: (1) the conventional thin-plate regression splines (TPRS) orthogonalised over the set of locations which exhibit an inherent ordering of smoothness, and (2) the cartesian product of two one-dimensional Fourier bases, which lack such canonical ordering. As a benchmark, we compare our mode-based estimator to the standard basis adjustment approach in regression, which is known to yield consistent results when basis functions are ordered by increasing roughness.
Figure 8: Results for the COVID-19 Data: Estimated mortality-rate ratio for an increase of $1\mu g/m^{3}$ increase in PM2.5 for different numbers of basis functions. The results are for two sets of basis functions: (a) when the basis functions have a canonical ordering of increasing roughness, and (b) when the basis functions do not have a canonical ordering among the basis functions. The blue line is for the model adjusted for $k$ basis functions along with the measured covariates, and the red is for basis voting, our mode-based estimator. The black dotted line is the estimate of 1.16 or 16% reported by guan2023spectral.
@@
As anticipated, when the basis set has an ordering of smoothness (as with TPRS), the results of both estimators are highly similar and align with the estimate from the previous analyses. Notably, when switching to the unordered Fourier basis, our mode-based estimator continues to produce estimates close to the presumed effect, whereas the standard basis adjustment method fails to control for unmeasured spatial confounding, producing a very different estimate. This result underscores the advantage of our approach in settings where the exposure does not represent a ‘rougher’ function of space than the confounder in terms of a given basis expansion.
@@ 7 Conclusion and Discussion
Unmeasured spatial confounding poses a major challenge for effect size estimation in spatial linear models, particularly when exposures and confounders share smooth spatial structure. Existing approaches typically assume exposures have some non-spatial noise, or at the very least, vary more roughly than confounders, but this assumption breaks down in many environmental health settings where exposures themselves are smooth or when using basis functions where there is no single notion of smoothness.
We address this gap by modeling exposures and confounders through basis expansions and show that the effect is identified under a plurality rule assumption on the basis functions in the support of the exposure but not the confounder. Leveraging this structure, we introduce basis voting, a plurality-rule based process to arrive at an estimator that identifies causal effects without requiring prior knowledge of valid candidate bases or bounds on the number of invalid ones. We provide theoretical guarantees for consistency and propose a provably more efficient candidate estimator with improved asymptotic variance. Simulations and a real data analysis confirm that our method delivers unbiased estimates whenever exposure and confounder signals are separable on a plurality of basis functions, regardless of their relative smoothnesses.
Plurality-rule based estimation provides a new, flexible and robust tool for addressing spatial confounding in complex applied settings, and this work advances spatial causal inference by relaxing the relative smoothness assumptions on which previous work relies. While our framework is limited to linear effects, we plan to explore future extensions to nonlinear models. Also, developing valid statistical inferential procedures using the mode-estimator is an important future direction.
@@ 8 Acknowledgement
The work was partly supported by grants from the National Institutes of Health and the Office of Naval Research. The authors acknowledge the use of ChatGPT.com for improving the exposition in some parts, code writing, and for some assistance with developing the theoretical results (some claims generated by the authors were proved with help from ChatGPT). All code and proofs were independently verified by the authors, who take full responsibility for the content. We also sincerely thank Professor Jonas Peters and Felix Schur for detailed and valuable feedback on an initial draft of the manuscript.
@@
S1 Comparison of Basis Voting and DecoR
We present a comparative evaluation of our proposed basis voting estimator against the DecoR method introduced by schur2025decor within a time-series framework. Although designed in a one-dimensional framework, DecoR is conceptually the closest to our basis voting framework. However, some aspects of the underlying philosophies behind the two methods differ substantially. In addition to contrasting their finite-sample behavior, we illustrate how basis voting naturally accommodates settings in which the number basis $d$ used in the analysis is not fixed, but instead allowed to increase with the sample size $n$. This flexibility is important as in practice, the number of basis functions used in the analysis typically grows with sample size.
DecoR formulates the confounding problem as an adversarial outlier problem after transforming the exposure and outcome into the frequency domain by projecting them onto all $n$ basis functions. Its performance crucially depends on the parameter $a$, which encodes prior knowledge about the proportion of transformed data points that must be treated as outliers—equivalently, the proportion of basis functions that generate invalid candidate estimates. The accuracy of DecoR therefore hinges on approximately correctly specifying $a$, making this choice critical in practice. Consistent with the modeling assumptions in schur2025decor, we use the correct basis specification for both DecoR and basis voting to ensure a fair comparison.
Figure 9 displays simulation results in a regime where the proportion of invalid basis-specific estimates increases with $n$. For our method, the mode is computed using all $n$ basis-specific regression estimates, directly highlighting that basis voting remains applicable even when $d=d(n)$ grows with sample size rather than remaining fixed. Across all simulations, the true parameter value is $\beta=2.5$. We consider three representative confounding scenarios:
  1. (a)
    Majority rule violation with homogeneous outliers.
    In this scenario, the majority of basis-specific estimates are contaminated, and all outliers yield the same biased value. As these outliers are majority, the mode is not at $\beta$ but at this biased value. So this is a scenario where both the majority and plurality conditions fail, and we therefore expect both DecoR and basis voting to be inconsistent. This case serves to illustratewhen either methods cannot recover $\beta$.
  2. (b)
    Majority rule holds with a fixed linear proportion of homogeneous outliers.
    A constant fraction of the basis functions produces the same incorrect estimate, but the uncontaminated basis functions (or valid candidates, in our terminology) still form a strict majority. Here, DecoR’s error increases with the error variance (as mentioned in schur2025decor), reflecting its sensitivity to noise and its reliance on the tuning parameter $a$. In contrast, plurality rule holds for this setting as the number of valid basis functions are the majority. We see this in the results where basis voting remains stable and demonstrates consistency across sample sizes.
  3. (c)
    Plurality (but not majority) rule holds with linearly growing outliers.
    For each $n$, a fixed proportion $p$ of the basis functions are outliers, and although they do not form a majority, their number grows linearly with $n$. The plurality rule holds while the majority rule fails. In this setting, basis voting is consistent whereas DecoR is not. This scenario demonstrates that the truncation threshold $d$ in Assumption 1 can grow with $n$, provided that the proportion of valid basis-specific estimates remains the largest group.
Figure 9: Basis-Voting vs DecoR performance across the three spatial confounding scenarios: (a) 60% bias of +2, 40% bias of 0, (b) 70% bias of 0, 30% bias of +2, (c) 35% bias of 0, 25% bias of +1, 20% bias of +2, 20% bias of +3, each evaluated under error variances $\sigma\in\{1,5,10\}$.
Together, these experiments highlight the distinct philosophies underlying the two approaches. DecoR requires specifying the proportion of outliers and casts confounding as an adversarial contamination problem, whereas basis voting does not rely on such prior knowledge and remains robust under both basis misspecification and growing number of basis functions used in the analysis as long as the plurality rule holds. These results underscore the flexibility and generality of the basis voting framework.
S2 Notations
@@
We denote $\mathbb{R}$ and $\mathbb{N}$ to be the set of real numbers and natural numbers, respectively. For a sequence of random variables, $\xrightarrow[n\to\infty]{P}$ denotes convergence in probability and $\xrightarrow[n\to\infty]{d}$ denotes weak convergence. We will also use stochastic ordering notations throughout the proof. For a sequence of random variable $\{X_{n}\}$ and a sequence of real numbers $\{a_{n}\}$, we say $X_{n}=\mathcal{O}_{P}(a_{n})$ if for all $\epsilon>0$ there exists a $M>0$, such that for all $n\in\mathbb{N}$, we have $\mathbb{P}\left(\left|\frac{X_{n}}{a_{n}}\right|>M\right)<\epsilon$. Also, we say $X_{n}=o_{P}(a_{n})$ if for all $\epsilon>0$, $\mathbb{P}\left(\left|\frac{X_{n}}{a_{n}}\right|>\epsilon\right)\xrightarrow[]{n\to\infty}0$. For a real valued function $f$ defined on an inner-product space $(\mathcal{M},\langle\cdot,\cdot\rangle)$, we denote $\|f(\cdot)\|^{2}_{\mathcal{M}}\coloneqq\sup_{t\in\mathcal{M}}\langle f(t),f(t)\rangle$. All the integrals are over the support $\mathcal{D}$ of the sampling density $f_{S}$, and hence unless specified, by $\int$ we will mean $\int_{\mathcal{D}}$. Recall that $\mathcal{F}(\mathcal{D},f_{S})$ is a space of continuous functions in $\mathcal{D}$ equipped with the inner product $\langle g,h\rangle_{f_{S}}\coloneqq\int_{\mathcal{D}}g(s)h(s)f_{S}(s)ds$ and the corresponding norm $\|\cdot\|_{f_{S}}$ with respect to the sampling measure $f_{S}$.
S3 Supporting Lemmas
Lemma S3.1 (Argmax Theorem (vaart1997weak)).
@@
Lemma S3.3.
[…]
where $||\cdot||_{2}$ is the spectral norm.
@@
$\bm{A}_{n}^{-1}=$ $(\bm{A}_{n}+\bm{A}-\bm{A})^{-1}$
$=$ $(\bm{A}+\bm{\Delta}_{n})^{-1}$
$=$ $\bm{A}^{-1}\left(\bm{I}+\bm{\Delta}_{n}\bm{A}^{-1}\right)^{-1}$
$=$ $\bm{A}^{-1}\sum_{k\geq 0}\left(-\bm{\Delta}_{n}\bm{A}^{-1}\right)^{k}$
[…]
This completes the proof of the proposition. ∎
Lemma S3.4.
If $K$ is Lipschitz continuous and the finite set $\bm{a}_{m}=\{a_{1},\cdots,a_{m}\}$ has a unique mode $t$, then for all $0<h\leq h_{0}$, where $h_{0}$ is some constant depending on $K$ and $\underset{a_{i}\neq t}{\min}|t-a_{i}|$, $\underset{x\in\mathbb{R}}{\operatorname*{arg\,max}}\ KD(\bm{a}_{m},K,h,x)=t$.
Proof.
and the uniqueness follows from $n_{t}>M$.
$\varepsilon\;<\;\frac{K(0)}{2m}\gamma.$ (8)
and thus
Evaluating at $x=t$ gives
$\begin{split}&KD(\bm{a}_{m},K,h,t)-KD(\bm{a}_{m},K,h,x)\\ \geq&\frac{1}{mh}\Bigl(n_{t}K(0)-(m-n_{t})\varepsilon-MK(0)-(m-M)\varepsilon\Bigr)\\ =&\frac{1}{mh}\Bigl((n_{t}-M)K(0)-(m+m-n_{t}-M)\varepsilon\Bigr)\\ \geq&\frac{1}{mh}\Bigl((n_{t}-M)K(0)-2m\varepsilon\Bigr)\\ =&\frac{2}{h}\Bigl(\gamma\frac{K(0)}{2m}-\varepsilon\Bigr)>0\end{split}$ (11)
By the choice (8), the right-hand side is strictly positive.
hence for any $x\in A$ and $0<h\leq h_{0}$,
Now fix any $r>0$, since
$KD(\bm{a}_{m},K,h,t)>KD(\bm{a}_{m},K,h,x),$
Hence $t$ is the unique global mode in this setup. ∎
Lemma S3.5.
Let $\{h_{k}\}_{k\geq 1}\subset\mathcal{F}(\mathcal{D},f_{S})$ an orthonormal system, fix $d\in\mathbb{N}$, assume that for each $j\leq d$, the zero set $\{s:h_{j}(s)=0\}$ has $f_{S}$-measure 0 (equivalently, $h_{j}^{2}>0$ $f_{S}$-a.e.). For $j\leq d$, set $V_{j}\coloneqq\operatorname{span}\{h_{k}:\,1\leq k\leq d,\ k\neq j\}$ and define the set $\mathbb{S}_{j}\coloneqq\{v\in V_{j}:\ \|v\|_{f_{S}}=1\}$ and $Q_{j}(v)\coloneqq\int h_{j}^{2}\,v^{2}\,f_{S}$. Then
$c_{\ast}\coloneqq\inf_{1\leq j\leq d}\ \inf_{v\in\mathbb{S}_{j}}Q_{j}(v)>0.$
Proof.
$Q_{j}(v)=\int h_{j}^{2}v^{2}\,f_{S}>0.$
Thus $Q_{j}$ is strictly positive on $\mathbb{S}_{j}$.
Lemma S3.6.
Let $\mathcal{H}=\{h_{j}\}_{j\geq 1}$ be orthonormal in $\mathcal{F}(\mathcal{D},f_{S})$, with $U=\sum_{j\geq 1}\alpha_{u}^{(j)}h_{j}$, $U_{\leq d}=\sum_{j\leq d}\alpha_{u}^{(j)}h_{j}$, $U_{>d}=\sum_{j>d}\alpha_{u}^{(j)}h_{j}$ for $d\in\mathbb{N}$ and define $S_{\mathrm{in}}\coloneqq\sum_{j\leq d}{\alpha_{u}^{(j)}}^{2}$. Assume:
  1. (i)
    $\max_{1\leq j\leq d}\|h_{j}\|_{\infty}\leq M<\infty$.
  2. (ii)
  3. (iii)
Then for every $j=1,\dots,d$ with $j\in O_{X,d}$ i.e. $\alpha_{u}^{(j)}=0$,
$\|h_{j}U\|^{2}_{f_{S}}>\|h_{j}U_{>d}\|^{2}_{f_{S}}.$
Proof.
We refer to Lemma S3.5 for this proof. Fix $1\leq j\leq d$ and assume $\alpha_{u}^{(j)}=0$ i.e. $j\in O_{X,d}$. Write $R_{j}\coloneqq\sum_{k\leq d,\,k\neq j}\alpha_{u}^{(j)}h_{k}$ and $T\coloneqq U_{>d}$. Set $A_{j}\coloneqq\int h_{j}^{2}R_{j}^{2}\,f_{S}\,ds,B_{j}\coloneqq\int h_{j}^{2}T^{2}\,f_{S}\,ds$ and $C_{j}\coloneqq\int h_{j}^{2}R_{j}T\,f_{S}\,ds.$ Then
By Cauchy-Schwarz, $|C_{j}|\leq\sqrt{A_{j}B_{j}}$, hence
Second, by Assumptions (i) and (iii),
Thus,
S4 Proofs of Theorems
S4.1 Proof of Theorem 1
Proof.
[…]
$\|KD(\mathcal{C}_{n,d},K,h,x)-KD(\beta+R_{d},X,h,x)\|_{\mathbb{R}}$
$=$ $\underset{x\in\mathbb{R}}{\sup}\left|KD(\mathcal{C}_{n,d},K,x)-KD(\beta+R_{d},K,x)\right|$
$=$ $\underset{x\in\mathbb{R}}{\sup}\left|\dfrac{1}{mh}\sum_{k\in S_{X,d}}K\left(\dfrac{x-\hat{\beta}_{k,n}}{h}\right)-\dfrac{1}{mh}\sum_{k\in S_{X,d}}K\left(\dfrac{x-(\beta+r_{k})}{h}\right)\right|$
$\leq$ $\underset{x\in\mathbb{R}}{\sup}\dfrac{1}{mh}\sum_{k\in S_{X,d}}\left|K\left(\dfrac{x-\hat{\beta}_{k,n}}{h}\right)-K\left(\dfrac{x-(\beta+r_{k})}{h}\right)\right|$
$\leq$ $\underset{x\in\mathbb{R}}{\sup}\dfrac{L_{K}}{mh}\sum_{k\in S_{X,d}}\left|\left(\dfrac{x-\hat{\beta}_{k,n}}{h}\right)-\left(\dfrac{x-(\beta+r_{k})}{h}\right)\right|\quad\text{(using Lipschitz Continuity)}$
$=$ $\dfrac{L_{K}}{mh^{2}}\sum_{k\in S_{X,d}}\left|\hat{\beta}_{k,n}-(\beta+r_{k})\right|\xrightarrow[n\to\infty]{P}0$
Now, based on the Plurality Rule Assumption 2 for some $d\geq d_{0}$, and Lemma S3.4, it is ensured that for $\theta_{0}=\beta$, we have $KD(\beta,K,h,\theta_{0})>\underset{\theta\notin G}{\sup}KD(\theta+R_{d},K,h,\theta)$ for any open set $G$ that contains $\theta_{0}$.
Thus defining $\hat{\beta}_{n}=\underset{x}{\operatorname*{arg\,max}}\ KD(\mathcal{C}_{n,d},K,h,x)$, we can say that $KD\left(\mathcal{C}_{n,d},K,h,\hat{\beta}_{n}\right)\geq\sup_{\theta}KD\left(\mathcal{C}_{n,d},K,h,\theta\right)-o_{P}(1)$. Thus, using Lemma S3.1, we can say that $\hat{\beta}_{n}\xrightarrow[n\to\infty]{P}\beta$. This completes the proof of the theorem. ∎
S4.2 Proof of Theorem 2
Proof.
Let us choose some $j\in\mathbb{N}$, such that $\alpha_{x}^{(j)}\neq 0$. Now, let us look at $\hat{\beta}_{j,n}^{PROJ}$:
@@
$\leq$ $\sum_{u=1}^{k}\left|\alpha_{x}^{(u)}\right|\mathbb{I}(u=j)\leq\left(\sum_{u\geq 1}\alpha_{x}^{(u)^{2}}\right)^{1/2}<\infty$
@@
Similarly we have $\frac{1}{n}\bm{h}_{n,j}^{\top}\bm{U}_{n}\xrightarrow[n\to\infty]{P}\alpha_{u}^{(j)}$. Lemma S3.2 implies that $\frac{1}{n}\bm{h}_{n,j}^{\top}\bm{\epsilon}_{n}\xrightarrow[n\to\infty]{P}0$. Thus when $\alpha_{x}^{(j)}\neq 0$, we have $\hat{\beta}_{j,n}^{PROJ}\xrightarrow[n\to\infty]{P}\beta+\alpha_{u}^{(j)}/\alpha_{x}^{(j)}=\beta+r_{j}$ which completes the proof of consistency. Let us define $A_{i}=h_{j}(S_{i})[U(S_{i})+\epsilon_{i}]$ and $B_{n}=h_{j}(S_{i})X(S_{i})$,which means $\frac{1}{n}\bm{h}_{n,j}^{\top}[\bm{U}_{n}+\bm{\epsilon}_{n}]=\frac{1}{n}\sum_{i=1}^{n}h_{j}(S_{i})[U(S_{i})+\epsilon_{i}]=\bar{A}_{n}$ and $\frac{1}{n}\bm{h}_{n,j}^{\top}\bm{X}_{n}=\frac{1}{n}\sum_{i=1}^{n}h_{j}(S_{i})X(S_{i})=\bar{B}_{n}$. We will try to derive the asymptotic distribution of $\bar{A}_{n}/\bar{B}_{n}$, which is the bias of the estimator $\hat{\beta}^{PROJ}_{j,n}$. Now we have seen that $\mathbb{E}A_{i}=\mathbb{E}h_{j}(S_{i})[U(S_{i})+\epsilon_{i}]=\mathbb{E}h_{j}(S_{i})U(S_{i})=\int h_{j}(s)U(s)f_{S}(s)ds=\alpha_{u}^{(j)}$ and $\mathbb{E}B_{i}=\mathbb{E}h_{j}(S_{i})X(S_{i})=\alpha_{x}^{(j)}$ Also, notice that
@@
Since $\int(h_{j}U(s))^{2}f_{S}ds<\infty$ and $\int(h_{j}X(s))^{2}f_{S}ds<\infty$ as the functions are elements of $\mathcal{F}(\mathcal{D},f_{S})$, by Multivariate Central Limit Theorem (CLT), we have:
@@
$\sqrt{n}\left(\dfrac{\bar{A}_{n}}{\bar{B}_{n}}-\underbrace{\dfrac{\alpha_{u}^{(j)}}{\alpha_{x}^{(j)}}}_{r_{j}}\right)\xrightarrow[n\to\infty]{d}N\left(0,\sigma^{2}_{j}\right)$ (15)